### v1

# Discretized Quantum Exhaustive Search for Variational Quantum Algorithms

Abstract
Quantum computers promise a great computational advantage over classical computers, yet currently available quantum devices have only a limited amount of qubits and a high level of noise, limiting the size of problems that can be solved accurately with those devices. Variational Quantum Algorithms (VQAs) have emerged as a leading strategy to address these limitations by optimizing cost functions based on measurement results of shallow-depth circuits. However, the optimization process usually suffers from severe trainability issues as a result of the exponentially large search space, mainly local minima and barren plateaus.
Here we propose a novel method that can improve variational quantum algorithms — “discretized quantum exhaustive search”. On classical computers, exhaustive search, also named brute force, solves small-size NP complete and NP hard problems. Exhaustive search and efficient partial exhaustive search help designing heuristics and exact algorithms for solving larger-size problems by finding easy subcases or good approximations. We adopt this method to the quantum domain, by relying on mutually unbiased bases for the $2^{n}$-dimensional Hilbert space. We define a discretized quantum exhaustive search that works well for small size problems. We provide an example of an efficient partial discretized quantum exhaustive search for larger-size problems, in order to extend classical tools to the quantum computing domain, for near future and far future goals. Our method enables obtaining intuition on NP-complete and NP-hard problems as well as on Quantum Merlin Arthur (QMA)-complete and QMA-hard problems.
We demonstrate our ideas in many simple cases, providing the energy landscape for various problems and presenting two types of energy curves via VQAs. Using our methods, we learn details that are similar to those obtained by experts’ knowledge, without using prior knowledge of the problem. We expect that merging prior knowledge and AI/ML tools in future work might improve far beyond current abilities of VQA.
1 Introduction
In the realm of modern computing, a new paradigm has emerged, promising a great computational advantage over classical computing. Quantum computing, a cutting-edge field at the intersection of physics, computer science, and engineering, harnesses the principles of quantum mechanics to perform computational tasks that are out of reach for classical computers.
The potential applications of quantum computing are vast and include solving complex optimization problems, simulating quantum systems, hacking modern cryptography, and revolutionizing artificial intelligence. However, this exciting promise also comes with substantial challenges, such as managing qubit stability, error correction, and hardware development.
Today’s quantum computers are often described as noisy intermediate-scale quantum (NISQ) computers due to the relatively low numbers of qubits available (e.g. 10’s to 100’s) and the relatively high levels of noise associated with them (e.g. decoherence and gate fidelity errors) [37, 15, 7]. These limitations result in quantum circuits with short width and shallow depth.
To make use of these NISQ machines, a class of hybrid quantum-classical algorithms that seek to leverage the relative strengths of quantum and classical computers is being developed. The most common example of such algorithms is Variational Quantum Algorithms (VQAs)[12]. VQAs use a quantum computer to prepare a short-depth Parameterized Quantum Circuit (PQC) representing a trial solution or ansatz to the problem at hand. Measurements of a final quantum state are used to calculate a cost function, which is then minimized on a classical computer to estimate the problem solution. Prominent examples of VQAs include the Variational Quantum Eigensolver (VQE) for quantum chemistry and materials applications [35, 24], Quantum Approximate Optimization Algorithm (QAOA) for combinatorial optimization problems [17], and Variational Quantum Linear Solver (VQLS) for linear algebra problems [9]. More details about these methods are provided in the appendix.
One of the main challenges associated with effective VQA implementations is related to the design of a suitable PQC/ansatz that balances expressibility and noise, while avoiding exponentially vanishing gradients of the cost function, referred to as the barren plateau (BP) problem [29].
Designing better ansatzes for VQAs is an active research area that has drawn much attention lately. There have been a number of attempts [44] to improve different ansatz approaches, including ADAPT-VQE [19] and Noise-Adaptive Search (QuantumNAS) [46].
Another important research area is developing methods for finding good initial state candidates for an ansatz. Finding a good initial guess for the optimizer is very important, as avoiding BP using smart ansatz engineering without drastically changing the problem’s search space might be impossible [11]. A good initial guess can make the difference between converging to the optimal solution or leading the optimizer to a local minimum trap. A good initial guess becomes more important as the problem size increases and the problem’s Hilbert space and the search spaces become exponentially larger. This increase in search size usually creates exponentially small gradients in the cost function landscape, making the optimization process almost impossible.
For many problems, there are already classical algorithms that can find an approximation to the optimal solution or wide knowledge about the problems at a small scale. These classical approximations, along with expert knowledge, can help create an initial guess for the ansatz for large scale problems. Still, there are problems where no good classical approximation can be found, and no useful prior knowledge is known about the problem. Therefore, some methods were previously proposed in order to find a good candidate for an initial guess for VQAs without the need to have prior knowledge.
We suggest to use a set of discrete points that are equally spread all over the search space of the problem. We define “discretized quantum exhaustive search” (DQES) in analogy to the use of standard exhaustive search in modern computer science.
As far as we know, no previous work has defined DQES or used such an approach to learn the cost function landscape when solving various problems, and as a set of initial states for VQA.
In addition, no method to sample the entire Hilbert space using evenly spaced states was defined. Using evenly spread states is important for gaining knowledge in a well-organized way about the landscape of solutions to a problem, and for drawing energy curves once applying optimizer for VQA, especially with the goal of bypassing barren plateaus.
Methods to create good initial states using random states were proposed [18, 31, 25]. Other methods proposed to use machine learning techniques [45, 38, 28]. Some ansatzes were proposed to handle barren plateaus instead of avoiding them [19, 20], using an adaptive ansatz structure (the ansatz changes during the optimization process, and can escape barren plateaus after entering one).
Another example of an initialization method that is suited for a general VQA is searching over states that can be created with only Clifford gates [39, 14]. Using circuits that contain only Clifford gates enables an efficient simulation of the quantum circuits on a classical computer, allowing one to search for an initial guess in a noiseless environment. Clifford states are spread evenly over the Hilbert space, which enables searching in different locations of the problem’s space and not only around specific locations. There are far too many Clifford gates in order to use them for a DQES approach (sampling an exponential, but reasonable number of states from the Hilbert space of a problem), as there are $2^{n^{2}+2n}\prod_{j=1}^{n}(4^{j}-1)$ Clifford states, where $n$ in the number of qubits [33]. $(2^{n})^{2}$ is much smaller than $2^{n^{2}}$, for example, for n=10, $2^{20}$ might be a reasonable number of samples, but $2^{100}$ becomes intractable.
The Clifford states have many benefits, but they are not suitable for every cost function, as the Clifford gates are stabilized by the Pauli operators. The expectation value of a Clifford state against a Pauli operator can lead to either an eigenvalue of $+1$ or $-1$, in case this Clifford is indeed stabilized by that particular Pauli operator, or to 00. For each Pauli operator, there are exponentially more Clifford states whose expectation value is 00 compared to those whose expectation value is either $+1$ or $-1$. Many common VQAs use Pauli expectation values to evaluate their cost function, as it can be implemented very efficiently and easily on a quantum computer.
The main goal of this research is to define novel methods to improve VQA ansatzes and initial states, leading to better accuracy and faster convergence rates. This goal is achieved by a few different novel methods -
  • Describing a new method to sample and explore VQAs’ problem space, using the defined DQES. This method allows one to understand the problem’s Hilbert space better, creating better guesses for initial parameter choices, and designing methods to avoid BP and local minima. This method is described in Sections 2.1 and 2.2. Examples of using this method are presented in Section 3.
  • Describing an alternative heuristic method to explore the VQA’s problem space, using an efficient approach. This method aims to bring a heuristical equivalent of the above method for larger problem instances while remaining computationally feasible. This method is presented in Section 2.4. Examples of using this method are also presented in Section 3.
1.1 Mutually Unbiased Bases
Mutually unbiased bases (MUBs) are orthonormal bases in a complex vector space that provide complementary and non-redundant information about the quantum states of different bases if measured [41, 47, 5, 16]. When two or more bases are mutually unbiased, it means that the measurement outcomes in one basis provide no information about the outcomes in another basis, and if you have a quantum state prepared in one basis and you measure it in another basis, the measurement outcomes are uncorrelated.
Definition 1.
For a finite-dimensional Hilbert space of dimension $d$, two bases $\Psi=\left\{|{\psi_{1}}\rangle,|{\psi_{2}}\rangle,\dots|{\psi_{d}}\rangle\right\}$ and $\Phi=\left\{|{\varphi_{1}}\rangle,|{\varphi_{2}}\rangle,\dots,|{\varphi_{d}}\rangle\right\}$ are considered mutually unbiased if the bases are orthonormal and the absolute value of the inner product between any vector from the first basis and any vector from the second basis is equal to $\frac{1}{\sqrt{d}}$.
$|\langle\psi_{i}|\varphi_{j}\rangle|=\frac{1}{\sqrt{d}},\textrm{ for every }i,j\in\{1,2\dots d\}$ (1)
Definition 2.
A set $\{B_{1},B_{2},\dots,B_{m}\}$ of orthonormal bases is called a set of mutually unbiased bases (a set of MUBs) if each pair of bases $B_{i}$ and $B_{j}$ is mutually unbiased.
It has been proven that for a Hilbert space with dimension $N$, the size of the largest possible set of MUBs is $N+1$. We call a set of $N+1$ MUBs a full set. For any $N$ which is a prime power ($p^{n}$, for a prime number $p$ and a natural number $n$), it has been shown that a set with such size exists [5, 47, 26, 16].
MUBs have applications in various areas of quantum information theory, which use the special properties of the mutual information between the bases. Some applications use MUBs to maximize the information gained from different measurements, for example in tomography experiments, like state tomography [47, 22, 1] and process tomography [32]. Other applications use MUBs to minimize the possible information gained by a third party during quantum communication in order to improve quantum key distribution and quantum cryptography [13, 36]. This is done by using the property of MUBs that the outcome is random when a measurement is made in a basis unbiased to that in which the sent state was prepared. When two remote parties share two non-orthogonal quantum states, attempts by an eavesdropper to distinguish between these by measurements will affect the system and this can be detected.
2 Methods
VQAs search for the optimal solution of the given problem using the search space imposed by the chosen ansatz. The Hilbert space of the problem is determined only by the problem definition and encoding and is not subject to changes by replacing the ansatz. The dimension of this Hilbert space is exponential in the number of qubits required to encode the problem into a quantum computer. Finding the optimal solution, or even a good approximation to it in such a huge space is usually a very hard task. In many search spaces, the gradients of the cost function become exponentially small in most areas of the search space, a phenomenon called barren plateaus (BP), and most of the states that have significant gradients are concentrated in a small area [29, 3]. The concentration of most of the cost function values that have a significant difference between them in a small area makes the search hard also for gradient-free optimizers [2].
Some ansatz designs try to narrow this problem by using an adaptive search scheme and by using prior knowledge of the problem to create a relatively small search space inside the exponential problem space [20]. Such prior knowledge is not always available, and even in the cases it does - reducing the search space usually comes with a cost of deeper quantum circuits, which results in more noise.
In our research, we propose methods to learn and gain insights about the values and structure of the cost function in the problem’s Hilbert space, in a way that is agnostic to the ansatz that will be used to try to solve the problem. Using such methods can help researchers design ansatzes, optimization algorithms, and initial guesses that can avoid BPs.
The core of these methods is to evaluate the cost function using predefined states that sample the Hilbert space to gain significant knowledge about the cost function landscape.
2.1 Discretized Quantum Exhaustive Search
Exhaustive search, also named ’brute force’, can solve small-size instances of NP-complete and NP-hard problems. By running exhaustive search on many small samples of a problem, as part of the research process, you can try and build heuristics on the general problem.
Here we define DQES which is useful for small size problems, and efficient partial DQES (for large-size problems), in order to extend these classical tools into the quantum computing domain, for near future and far future goals. We do so by relying on mutually unbiased bases for $2^{n}$-size Hilbert space. For non-partial DQES, we limit ourselves to small $n$.
Our method enables obtaining intuition on NP-complete and NP-hard problems as well as on Quantum Merlin Arthur (QMA)-complete and QMA-hard problems.
NP problems can be solved in polynomial time by a classical computer once a relevant and correct hint is given from an external source. In a similar way, the QMA complexity class contains problems that can be solved by a quantum computer in polynomial time once a relevant hint is given by an external source. We notice that the method of VQE resembles the idea of using a hint, yet instead of obtaining the correct hint as in QMA problems, clever scientists guess a sequencial set of hints. When a quantum computer tries to solve an NP problem of a small size, it is possible to test all possible classical hints via the exhaustive search, to gain knowledge on the structure of the problem, and design heurstics for larger size problems.
As looking over all the possible states as possible hints is not possible — for a quantum space — we look for an approximated method relying on evenly spread hints, that can resemble the classical exhaustive search.
Searching an exponential number of different states, which are spread evenly all over the problem’s relevant Hilbert space, for small size problems, may help designing heuristics that can then be used for larger-size problems. Specifically, it may help studying the energy landscape of problems hence the locations of barren plateaus and local minima.
2.2 MUBs for Exploring The Cost Function Landscape
The MUB states of a full set of MUBs have unique properties that make them good for extracting information for purposes such as state tomography, and process tomography. In a similar manner, we suggest that these states can be used to extract information about the landscape of a cost function of a given problem, if used accordingly. A full set of MUB states can be a good candidate as states for a DQES for small-size problems. There is an exponential number of states that are spread all over the Hilbert space, and their information is, in some sense, mutually exclusive.
By measuring the cost function value of a given problem at all these states, one can get evenly-spread (through space) information about the structure of the cost function. Such information can be properties of states that give similar values, hinting that there could be barren plateaus areas in the cost function landscape, properties of states that potentially present an area with local minima, and properties of states that have a larger variety of results, pointing to areas that might be interesting to search for the optimal result.
2.3 Circuit Representation of MUB States
The computation space of a quantum device consisting of $n$ qubits is of dimension $2^{n}$. It has been shown [5, 47, 26, 16] that for such a space there exists a set of $12^{n}+$ 1 orthonormal bases which are mutually unbiased. Finding such a set of mutually unbiased bases for any number of qubits is not a trivial task. Few methods have been proposed to find a full set of unbiased bases, containing $12^{n}+$ 1 orthonormal bases [27, 6, 5, 47]. As far as we could figure out, the computation cost of any such known algorithm to find orthonormal bases which are part of a full MUB set, is exponential in the number of qubits. For small $n$ we noticed one research that found efficient ways to implement these MUBs in the lab using adjusted measurement equipment [21]. We need to have a quantum circuit that prepares each state in those bases. Known algorithms to find a quantum circuit that creates an arbitrary given quantum state take exponential time to compute in the worst case and might produce a very deep quantum circuit, therefore usually can not be used in the VQA context.
Here we investigate two different circuit representations of MUB states, both are sufficiently shallow for our demonstrations. In the first representation, each basis in the MUB set is defined by a quantum circuit that transforms a state in the computational basis to a state in that basis. We have manually found such shallow circuits for each basis in a full set of MUBs of 2 and 3 qubits, see Fig.1. This method provides only a discrete map, yet it may be used for finding an optimal initial state (or a set of states), that can then be used as the initial state(s) of a parametrized ansatz controlled by an optimizer as in all usages of VQA.
In contrast, the second representation relies directly on a chosen ansatz circuit. In this representation, each MUB state is defined by the parameters vector of an ansatz which prepares the state, hence each MUB state can then directly be used by utilizing the optimizer to modify the parameters. When using that method, the chosen ansatz must be expressive enough to allow the preparation of each of the states of the MUBs. In Fig. 1, subfigure (c), we give a hardware-efficient ansatz circuit for generating 2-qubit MUBs. When acting on the computaitonal state $|{i}\rangle$, it gives the i’th element of the MUB basis. Note that the circuit in 1 (c) can be used to generate all MUB states directly by adding a single layer of NOT gates (where needed) at the start of the circuit, if the initial state is always the all-zeros state.
Figure 1: Circuits transforming states from the computational basis to each basis in (a) a 2 qubit MUB set. (b) 3 qubit MUB set. (c) 2 qubit MUB set and the relevant parameters of a hardware efficient ansatz to create these circuits.
2.4 Efficient Partial Discretized Quantum Exhaustive Search
Another tool we suggest and demonstrate via simple examples is an efficient partial DQES. Such a search may also provide heuristics and potentially lead to exact algorithms, for solving larger-size problems, by finding easy subcases or better than classical approximations.
Relative to existing VQA methods that rely on a single initial state or just a few initial states, the method we suggest relies on a polynomial number of initial states. To obtain a scalable method we let some qubits be much more dominant than others in the cost function of a VQA problem, yet we do not assume that the dominance of each qubit is known a-priori. We suggest a method to search possible initial states by using a specific set of partial-MUB states.
Our chosen partial MUB set is defined as follows. Let $n$ denote the number of qubits in the problem. Let $K$ be a constant which we call the “partial MUB constant”. A partial-MUB state is a tensor product between $n-K$ qubits that form any single state (e.g., WLOG, the single state may be the all-zero state in the computational basis), and $K$ qubits that form a state that is part of one of the bases in a full MUB set. For small values of $K$ we may generate all MUB states. For the scalablity proof we shall not assume tha $K$ is small.
For a constant $K$, The number of different options to choose the qubits splitting is (${{n}\choose{K}}=O\left(n^{K}\right)$. The number of MUB states for $K$ qubits is $(2^{K}+1)\cdot 2^{K}=(2^{K})^{2}+2^{K}$, leaving us with a total number of
(${{n}\choose{K}}\cdot((2^{K})^{2}+2^{K})=O(n^{K}\cdot 4^{K})=O(n^{K})$ (2)
states.
When choosing $K=\log(n)$, the number of MUB state options becomes $n^{2}+n$, and the number of different qubit splitting choices is (${{n}\choose{\log(n)}}=n^{\log(n)+O(\log(\log(n)))}$. The result is super-polynomial but still sub-exponential.
2.5 MUB States as an Initial Guess for Ansatzes and Optimizers
One possible outcome from the sampling of the cost function can be a good initial guess or a set of such initial guesses to use for the optimization process, which might help avoid barren plateaus and/or local minima and increase the chances of converging into the global minimum. The choice for the state to use as an initial guess can simply be the state that gave the minimum value in the cost function evaluation (as long as no additional prior knowledge is known about the problem or the structure of the optimal state), or the best state among a basis whose states showed large differences between their cost values, hinting for an area within the cost function landscape with high gradients.
To the best of our knowledge, the existing gate-based ansatzes’ circuits do not parameterize all of the quantum gates inside them, and all problem-agnostic ansatzes do not have a set of parameters that set the circuit to be equivalent to the identity. As a result, if a circuit that prepares a desired initial guess is inserted at the start of the ansatz, the first sampling point of the optimization will be different than the desired initial guess, even when the zero parameters vector is used. In order to overcome this, we suggest two methods -
  • Expressive ansatz - Use an expressive ansatz, that can generate (given the relevant parameters vector) each of the MUB states that were used for the energy landscape sampling. Then, when wishing to perform an optimization algorithm and use a MUB state $|\psi\rangle$ as an initial guess, use the parameter vector that makes the ansatz generate $|\psi\rangle$ as the initial guess for the optimization. Examples of using such expressive ansatzes to represent MUB states were presented in section 2.3.
  • shifted MUBs - Choose an ansatz before sampling the energy landscape, let the ansatz be $U(\theta)$. Choose a fixed parameters vector $\theta_{0}$ (for example, the all-zero parameters vector). Applying the unitary operation $U(\theta_{0})$ to a set of MUB states preserves their mutually unbiased properties (states from each basis will remain orthogonal, and the inner product of states from different bases will remain the same). Therefore, instead of sampling the energy landscape using MUB states $|{\psi_{1}}\rangle,|{\psi_{2}}\rangle\dots$, one can use the shifted MUB states $U(\theta_{0})|{\psi_{1}}\rangle,\;U(\theta_{0})|{\psi_{2}}\rangle\dots$ Then, in order to use $U(\theta_{0})|{\psi_{k}}\rangle$ as an initial guess for the optimization, one can use the circuit preparing the MUB state $|{\psi_{k}}\rangle$ followed by the ansatz with the fixed parameters vector $\theta_{0}$.
As many research papers show [18, 45, 28, 38], choosing a good initial guess (or guesses) for the optimization process of a VQA might affect dramatically the convergence of the algorithm, namely, might affect either (or both) the number of iterations required for convergence and the quality of the final result to which the algorithm converges.
For some problems, there are domain-specific initial guesses, which use prior knowledge of the problem. For example, a common initial guess for finding the ground energy of different molecules, which is found using the Variational Quantum Eigensolver, is the Hartree-Fock state (note that finding it is an NP-complete problem). Such initial guesses do not always produce the best results, and are only available for problems in which a prior knowledge of the problem is present.
In this subsection, we suggested a method to search for an initial guess that can be applied to any VQA, without the need for prior knowledge, which is useful for small-size problems. Of course, for larger sizes, prior knowledge or AI/ML tools are highly useful. Then, by using partial MUB states, one can sample a polynomial number of states, in order to find a state which has a better chance to lead to a convergence to an optimal result.
3 Results
We used the method of DQES on various problems. For problems that require up to 3 qubits to encode the cost function, we used all the states of a full MUB set. Calculating the cost function at all of the states of a full MUB set allows us to sample evenly from the entire problem’s landscape. As the number of states in a full MUB set grows exponentially in the number of qubits, for problems that require more than 3 qubits to encode the cost function we used the partial DQES method, by using either 2-qubit MUB states or 3-qubit MUB states and keeping the rest of the qubits of the problem in the zero state. For each MUB state, we have calculated the cost function for this state analytically without the need to measure multiple copies of the state (therefore, the result does not have statistical errors). Such analytical calculation is possible when one has an explicit and efficient representation of the quantum state and the number of non-identity terms in the operators of the cost function is small. When using the partial DQES, for each state we have repeated the calculation for each possible allocation of the MUB qubits as part total qubits number. For example, when using a 3 qubits MUB state for a problem with a size of 4 qubits, we repeated the process 4 times for each 3-qubit MUB state - each time another qubit was kept in the $|0\rangle$ state, while the 3 other qubits were assigned as the MUB state.
For some of the problems, we used the best resulting MUB states as an initial guess for the VQE algorithm. When running the VQE algorithm, having an explicit representation of the quantum state requires exponential (in the number of qubits) memory, therefore calculating the result analytically is possible only for small size problems. The order of the bases in the plots in this section is the same as listed in Chapter 2.3, in Figure 1 (a) and (b). The order of the states inside each basis is by the computational basis states - $|00\rangle,|01\rangle,|10\rangle,|11\rangle$ for the 2 qubits MUB states and $|000\rangle,|001\rangle,|010\rangle,|011\rangle,|100\rangle,|101\rangle,|110\rangle,|111\rangle$ for the 3 qubits MUB states. Each state in the other bases can be defined as the state resulting from evolving one of the computational basis states using the basis transformation matrix (or the quantum circuit representing the basis). The order of the states is defined by the states that are produced from the states in the computational basis in their regular order, transformed to the other basis in this manner.
3.1 Single qubit example
In order to provide intuition regarding the use of MUB-states in VQA, we make use of the simplest case — the case of a single qubit. In this extremely simple case, we can provide a geometrical interpretation of the DQES search and the optimization process that follows it. Unfortunately, we are not aware of a possibility to provide such a geometrical intuition for larger dimentsions.
For a single qubit, the canonical MUBs are the $X$, $Y$, and $Z$ bases, and of course these could be transformed via an arbitrary rotation to any other, non-canonical, shifted MUB.
Among all VQAs, not many can be applicable to a single qubit. However, calculating an eigenvector is applicable: For our demonstration, we choose the Hermitian matrix $\sigma_{x}+\sigma_{y}$, and the goal is to find its eigenstate. The ansatz has three parameters since an arbitrary single-qubit rotation is defined by the three Euler angles.
We started with each of the six MUB-states, generated using an ansatz, and we applied the COBYLA optimizer. For each initial state, the first guess is the MUB state, and the next three guesses are provided (with no optimization attempt) by COBYLA to be a modification of a single angle of the three parametrized angles, by a single pre-defined “distance”. Then, starting from iteration number 4, the COBYLA algorithm chooses the next guess. There are two different ways to stop the iterations: either by the optimizer finding no way to improve the current cost value, or by the cost function reaching a pre-defined threshold. We present two examples of reaching the solution — starting from two different MUB states, as well as a closeup on the path in the Hilbert space (in this case, the Bloch sphere) at the last iterations. In figure Fig.2 we see those optimization paths examples.
Figure 2: Optimization paths (energy curves) for 1 qubit search space example. (a) Optimization path obtained by COBYLA for a single-qubit example of the use of MUB states for VQA: the full path, starting from state |ket|$\rangle$| - ⟩. (b) Optimization path obtained by COBYLA for a single-qubit example of the use of MUB states for VQA: the full path, starting from the state $|-_{i}\rangle$. (c) A closeup on reaching the solution for the optimization path presented in (a). (d) A closeup on reaching the solution for the optimization path presented in (b).
3.2 Molecular electronic structure problems
We have calculated the energy of various Hamiltonians, and used a full set of MUB states of 2 qubits for the partial DQES method. In all of these Hamiltonians we used the Born-Oppenheimer approximation and used STO-3G basis as the basis for the orbitals function approximation (solving the FCI problem). The Hamiltonians were converted into spin Hamiltonian using parity transformation [42], and encoded into qubits after utilizing symmetries in the Hamiltonian to reduce the amount of qubits needed in order to encode the solution by 2 (as described in [10, 44]). All of the Hamiltonian coefficient calculations were done using the “qiskit-nature” Python package[40], version 0.4.3. When we have calculated the Hartree-Fock states for comparisons, we used the same parameters (STO-3G basis as basis for the orbitals and the same atomic distance) and used the “pyscf” Python package[43], version 1.6.3, to do the calculation. Some preliminary details are provided in the appendix.
The first Hamiltonian represents the $H_{2}$ molecule at an atomic distance of $0.75[A]$. 2 qubits were required to represent this Hamiltonian, as the relevant orbitals included in the calculation are the 1s orbitals from each $H$ atom, resulting in 4 spin orbitals. By utilizing symmetries we managed to reduce the number of qubits required in the encoding to 2. The resulting Hamiltonian, in the tensored Pauli operators form, is:
$\begin{split}&\mathbb{H}_{H_{2}}^{r=0.75}=-1.05540303\cdot II+0.38874759\cdot IZ-0.38874759\cdot ZI\\ &-0.01117714\cdot ZZ+0.18177154\cdot XX\\ \end{split}$ (3)
The results are shown in Fig.3. The best MUB state result is the same as the Hartree-Fock state energy. By using the Real Amplitudes Hardware Efficient Ansatz with a single layer, and using the 3 best MUB result as an initial state, we were able to achieve final results within chemical accuracy (within $1.6[mH]$ from the FCI results) in all 3 cases. The ansatz is expressive enough to create all the MUB states, so we used the parameters form of the MUB states representation to initialize the ansatz with these states, as shown in Fig. 1 (c). We achieved these results using a noiseless simulator and the COBYLA optimizer. The convergence results (energy curves — change of energy as a function of the iterations in the optimization process) are shown in Fig.4. All 3 cases converged to approximately the same result.
Figure 3: Exhaustive search of all the 2 qubit MUB states and their energy respect to the $H_{2}$ molecule in 0.75 [A] atomic distance.
Figure 4: Convergence results of the optimization process of a VQE algorithm on the $H_{2}$ molecule in 0.75 [A] atomic distance using a noiseless simulator, the COBYLA optimizer, and different initial states. The initial states are the 3 MUB states that had the lowest energy in the landscaping of the molecule. The right figure is a zoom to the final iterations of the optimization.
The second Hamiltonian represents the $HeH^{+}$ molecule at an atomic distance of $1[A]$. 2 qubits were required to represent this Hamiltonian, as the relevant orbitals included in the calculation are the 1s orbitals from both the $H$ atom and the $He$ atom, resulting in 4 spin orbitals. By utilizing symmetries we managed to reduce the number of qubits required in the encoding to 2. The resulting Hamiltonian, in the tensored Pauli operators form, is:
$\begin{split}&\mathbb{H}_{HeH^{+}}^{r=1}=-3.04506092\cdot II+0.50258052\cdot IZ+0.11926278\cdot IX\\ &-0.50258052\cdot ZI+0.11926278\cdot XI-0.13894646\cdot ZZ\\ &-0.11926145\cdot ZX+0.11926145\cdot XZ+0.11714671\cdot XX\\ \end{split}$ (4)
The results are shown in Fig.5. The best MUB state result is the same as the Hartree-Fock state energy. By using the By using the Real Amplitudes Hardware Efficient Ansatz with a single layer, and using the 3 best MUB result as an initial state, we were able to achieve final results within chemical accuracy also for this problem. We achieved these results using the same simulator and optimizer. The convergence results are shown in Fig.6. Although all 3 cases converged to an energy within chemical accuracy, there is a slight energy difference between the converged energies. The state that had the lowest energy in the landscaping ($|01\rangle$) converged to the lowest energy in the optimization process.
Figure 5: Exhaustive search of all the 2 qubit MUB states and their energy respect to the $HeH^{+}$ molecule in 1 [A] atomic distance.
Figure 6: The convergence results (energy curves — change of energy as a function of the iterations) of the optimization process of a VQE algorithm on the $HeH^{+}$ molecule in 1 [A] atomic distance using a noiseless simulator, the COBYLA optimizer, and different initial states. The initial states are the 3 MUB states that had the lowest energy in the landscaping of the molecule. The right figure is a zoom to the final iterations of the optimization.
The third Hamiltonian represents the $LiH$ molecule at an atomic distance of $1.5[A]$. 10 qubits were required to represent this Hamiltonian when taking into consideration all of the relevant orbitals (6 orbitals, which represent 12 spin orbitals). We used the partial DQES method, defined in 2.4. We used 2 qubits MUB states, and kept the rest of the qubits at the zero state. For each experiment, we tested all the possible pairs of qubits to be defined as the qubits of the MUB state. We tested how the results vary when reducing the amount of orbitals which we took into consideration in the encoding. Each removed orbital reduces the amount of needed qubits in the encoding by 2. The FCI result shown in the figures resembles the FCI result of the tested subspace, meaning - when not considering the removed orbitals. The FCI energy result is different when considering more orbitals and different basis functions, as shown in [4].
The results when using 10, 8, 6, 4 and 2 qubits in the encoding (using orbitals [0,1,2,3,4,5], [0,1,2,3,5], [0,1,2,5], [1,2,5], [1,2] respectively) are shown in figure Fig.7. The best MUB state result is the same as the Hartree-Fock state energy in all of those cases.
Figure 7: Exhaustive search of all the 2 qubit MUB states and their energy respect to the $LiH$ molecule in 1.5 [A] atomic distance when considering different number of orbitals in the calculation. (a) considering 2 orbitals in the calculation. 2 qubits are required in this encoding. (b) considering 3 orbitals in the calculation. 4 qubits are required in this encoding. (c) considering 4 orbitals in the calculation. 6 qubits are required in this encoding. (d) considering 5 orbitals in the calculation. 8 qubits are required in this encoding. (e) considering 6 orbitals in the calculation. 10 qubits are required in this encoding.
From these results, it can be seen that the orbitals whose occupation is the most relevant for the optimal result are orbitals 1 and 2. In addition, it can be seen that the most relevant information is found in the computational basis. Both of these results are known when using the prior knowledge of quantum chemistry. Such information might be gained also for problems where no such prior information is found, like some cases of a system of linear equations, which can be solved using the VQLS algorithm. By gaining such information, an initial guess can be made for the algorithm, and a better choice of ansatz can be made (for example - choosing an ansatz that does not use phase gates at all, like the Real Amplitudes ansatz). The results of the LiH experiments are summarized in Fig.8.
Figure 8: Best exhaustive search results for different amounts of qubits used in the encoding compared to the Hartree-Fock energy and the subspace FCI results.
3.3 Transverse-Field Ising Problem
We used MUB states to sample the energy landscape of transverse-field Ising Hamiltonians with random coefficients for the $ZZ$ and $X$ terms. The Transverse-Field Ising model Hamiltonian is
$H=\sum_{i,j}c_{ZZ}Z_{i}Z_{j}+c_{X}X_{i}$ (5)
while $c_{ZZ}$ and $c_{X}$ are the coefficients of the $ZZ$ and $X$ terms respectively, which were chosen randomly in our tests.
In Figures  9 and  10, we show results of energy landscaping of such Hamiltonians with 3 qubits. The landscaping is done using a full MUB set. We colored the results from different bases and show that by having different parameters for the Hamiltonian, the basis that has a state with the lowest energy is different. Because of the locality of each term in the sum of operators that create the Hamiltonian, MUB states from high dimension can be traced out to dimension 2 and their expectation value can be calculated exactly using a classical computer. Such calculation can be done only if the state is known and can be represented directly as a quantum state and not only as a unitary transformation that evolves some initial state (which in that case such transformation should be calculated first, and in case of a very high order, might not be possible to calculate on a classical computer). For that reason, such a calculation can’t be used during the optimization process after using a MUB state as the initial state for an ansatz.
Figure 9: Energy landscape of a transverse-field Ising Hamiltonian. each point represents a different state, while all of the states from the same basis have the same color. The lowest energy state is part of the computational basis (blue color). The Hamiltonian coefficients are: $H=0.04645122\cdot ZZI+0.04645122\cdot IZZ+0.27498273\cdot XII+0.27498273\cdot IXI+0.27498273\cdot IIX$
Figure 10: Energy landscape of a transverse-field Ising Hamiltonian. each point represents a different state, while all of the states from the same basis have the same color. The lowest energy state is part of the Hadamard basis (orange color). The Hamiltonian coefficients are: $H=0.61436456\cdot ZZI+0.61436456\cdot IZZ+0.32435029\cdot XII+0.32435029\cdot IXI+0.32435029\cdot IIX$
3.4 Combinatorial Problems
We used a full set of MUB states of 3 qubits for the efficient partial DQES method to explore the landscape of the Max-Cut problem (additional details in the appendix). We used 3 qubits MUB states, and kept the rest of the qubits at the zero state. For each experiment, we tested all possible triples of qubit locations to be defined as the qubits of the MUB state.
As an example, we used a random graph with 8 nodes, and sampled it’s energy landscape using partial-MUB states. We can see from the spread of the different results, that the cost function has the most differences in values in the computational basis, while in the other bases, most of the states yield the same result. Such knowledge can be used to limit the ansatz to specific states in the Hilbert space of the problem. We also tried to find the optimal solution by running an optimization process using the VQE algorithm (using COBYLA as optimization algorithm). We compared the results between different initial state - 2 MUB states which had low energy in the energy landscaping step, and 2 random states which had low energy relative to other random states that were randomized (we sampled the same amount of random states and MUB states). The results are shown in Fig.11
Figure 11: Max-Cut problem example. (a) Exhaustive search of all the possible tensor products of 3 qubit MUB states with the rest of the qubits in the zero state, and their cost function value for a random 8 nodes Max-Cut problem. (b) Convergence graphs of VQE runs on the 8 nodes Max-Cut problem, starting with different initial guesses - 2 choices of MUB states and 2 random states.
4 Conclusion
In this paper, we have tried to tackle the most dominant problems, local minima and barren plateaus, of one of the promising algorithmic approaches to the NISQ era of quantum computing - VQAs.
We have developed a method to efficiently sample the Hilbert space of the cost function of a VQA problem in a efficient partial DQES approach. We used MUB states in order to extract maximum information out from such sampling, in an ansatz agnostic method. We tested our method on multiple different problems, including different molecular electronic structure problems, Transverse-Field Ising model problems, and Max-Cut problems.
We showed that such a method can be used in order to learn characteristics of the landscape of the cost function and some useful information about the given problem at hand, like good approximations of the problem and good initial guesses to the solution. We managed to learn good approximations and initial guesses regarding the tested problems, which are similar to those achieved by experts in the problems’ fields. We have extracted such information without using prior expert knowledge. We believe that the information learned on the cost function landscape can be used to help avoid barren plateaus during the optimization process done in a VQA.
4.1 Open problems for Future research
Our research introduces new ansatz and methods to deal with the trainability of VQAs, but further research is still needed in order to make these algorithms achieve results better than the best classical computers using today’s available quantum hardware. Such further research might include -
  • Improve DQES and efficient partial DQES and try these on additional problem types.
  • Try to use the information gained from the DQES search or the efficient partial DQES to avoid BP and improve the algorithm convergence rate or final accuracy.
  • Apply DQES and partial DQES to problems with larger number of qubits.
  • Merge our methods with experts’ prior knowledge of specific problems.
  • Merge our method with various AI and ML tools.
  • Check our methods with various optimizers, and define novel types of optimizers to deal with the set of MUB states or a set of partial MUB states.
Acknowledgements:
D.M. and T.M. thank the Helen Diller Quantum Center at the Technion for their generous Support. T.M. and I.A. thanks the Quantum Computing Consortium of Israel Innovation Authority for financial support.